# Hypergraph Representations of scRNA-seq Data for Improved Clustering with Random Walks

### v3

@@ III.4.1 Evaluation Metrics
Dataset Greedy Modularity Louvain Infomap Leiden PCA t-SNE Node2Vec UMAP EDVW DIPHW CoMem CoMem-DIPHW
scMixology ARI 0.5836 0.1976 0.2327 0.9933 0.9900 0.9898 0.9900 0.9898 0.6619 0.9868 0.9933 0.9900
scMixology NMI 0.7193 0.3818 0.3843 0.9864 0.9797 0.9810 0.9797 0.9810 0.6609 0.9743 0.9864 0.9797
scMixology 5 class ARI 0.6645 0.3374 0.2998 0.6746 0.9873 0.9867 0.9461 0.9838 0.6873 0.9373 0.9880 0.9923
scMixology 5 class NMI 0.7852 0.4498 0.4355 0.7917 0.9747 0.9748 0.9148 0.9707 0.7740 0.9001 0.9770 0.9849
Table 3: Comparison of ARI and NMI across Clustering Methods for the ScMixology [61] Benchmark Datasets. The table is divided into 5-class (top) and 3-class (bottom) results. Best values are in bold, second-best are underlined. While PCA also performs well on these simple clustering tasks, CoMem and CoMem-DIPHW show robust performance across both datasets and are expected to outperform on more complex, weakly modular data, based on the previous simulation results.
III.6 Human Pancreas Results
@@
For datasets without ground-truth annotations—human pancreas, mouse pancreas, human brain, and mouse brain—we conducted differential expression (DE) analysis based on clustering results. We present our differential expression analysis results from the human pancreas dataset by CoMem-DIPHW and PCA in detail and include a brief overview of results from the other datasets in the Supplementary Materials Section  Supplementary Materials.
@@
We report the top 10 differentially expressed genes (DEGs) from each human pancreas cell cluster identified by CoMem-DIPHW and PCA in Fig. 7a and Fig. 8a respectively. To evaluate clustering performance, we examined the expression of these DEGs across clusters (Fig. 7b) and PCA (Fig. 8b). Violin plots show that the DEGs identified by CoMem-DIPHW exhibit more cluster-specific expression patterns than PCA-derived DEGs, which tend to show higher expression in multiple clusters. For example, the average expression of the DEGs from PCA cluster 1 is higher in clusters 2, 5, and 6. Similarly, the DEGs from cluster 7 show higher average expression in cluster 6. In general, the average expression of the DEGs from the PCA clusters is less cluster-specific than the DEGs from the CoMem-DIPHW clusters.
To tentatively label the cell clusters, we used canonical gene markers reported in the original human pancreas scRNA-seq study [58]. We assigned clusters based on the expression of these canonical markers including GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells). Each marker is distinctly expressed in a well-defined cluster in the embedding generated by CoMem-DIPHW, as shown in Fig. 7d, confirming that CoMem-DIPHW embeddings effectively distinguish different cell types. In contrast, PCA embeddings lacked clear separation between different cell types (Fig. 8d).
For example, in Fig. 7d, cluster 0 was identified as acinar cells by the high expression of PRSS1. Clusters 2 and 7 were identified as alpha cells by the high expression of GCG. Cluster 3 was classified as ductal cells by the high expression of KRT19. Both COL1A1 and ESAM are highly expressed in cluster 4, suggesting mesenchymal or endothelial cells. Cluster 5 was classified as delta cells by the high expression of SST, while cluster 6 was classified as beta cells by the high expression of INS. The specific DEGs supporting each classification are detailed in Fig. 7a.
@@
To further validate the cell type assignments and clustering results, we cross-referenced the cluster-specific DEGs (see Fig.7a) with markers from the PanglaoDB cell type marker database (version 27 Mar 2020)[74]. The assignment of cell types was determined based on the overlap between the DEGs of each cluster and the cell type-specific markers in the PanglaoDB marker database, with a match score computed based on the proportion of matched markers (Table 7c). These results align well with our previous analysis based on marker gene distribution (Figure 7d). The only discrepancy lies in cluster 4. Matching the DEGs with PanglaoDB cell type markers suggests Pancreatic stellate cells (PSCs) or Granulosa cells, while the COL1A1 and ESAM expression plots indicate mesenchymal or endothelial cells. However, this apparent discrepancy can be explained by the fact that PSCs can differentiate into cells with endothelial-like properties [75], and granulosa cells can exhibit mesenchymal-like behaviors [76].
@@
Figure 7: Clustering Performance of CoMem-DIPHW on the Human Pancreas Dataset and Cell Type Annotation using Differentially Expressed Genes (DEGs) and Canonical Markers. (a) Top 10 DEGs per cluster identified by CoMem-DIPHW. For example, Cluster 0 (acinar cells) includes genes associated with digestive enzyme production, a primary function of pancreatic acinar cells, including CPA1, CPA2 [68], PLA2G1B [69], CTRB1 [70], and PRSS1 [71]. Clusters 2 and 7 (alpha cells) include markers GLS, CRYBA2, FAP, LOXL4, MAFB, and RGS4 (cluster 2) and LOXL4, PAX6, RGS4, and CRYBA2 (cluster 7). Cluster 3 (ductal cells) includes SLC4A4, ANXA4, KRT19, CFTR, CLDN1, and ONECUT2. Cluster 4, likely mesenchymal or endothelial cells, aligns with 9 of the top 10 DEGs found by CoMem-DIPHW, including COL4A2, COL4A1, SPARC, EMP1, IGFBP4, HTRA1, PXDN, LAMA4, and NID1. Cluster 5 (delta cells) includes GAD2, PCSK1, LEPR, and CASR, while Cluster 6 (beta cells) includes MAFA, ERO1LB, PDX1, IAPP, SYT13, HADH, and ADCYAP1. Cell type assignments are based on canonical marker expression patterns shown in (d). (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high within-cluster expression of DEGs and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and PanglaoDB cell type-specific markers, with match scores computed based on the proportion of matched markers. (d) Expression of human pancreatic canonical markers in CoMem-DIPHW embeddings. UMAP visualizations show CoMem-DIPHW cell embeddings, with cluster IDs assigned by K-means clustering. Each subplot displays the expression of a canonical marker gene for a specific human pancreatic cell type. The cell-type-specific markers considered include: GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells). Color intensity represents gene expression levels, with red hues indicating higher expression. Clusters show high expression of cell-type-specific canonical markers are assigned to the corresponding cell types. Overall, clustering quality is demonstrated through (b) high within-cluster DEG expression, (c) consistency between PanglaoDB-based annotation and clustering, and (d) distinct separation of cell types.
@@
Figure 8: Clustering Performance of PCA on the Human Pancreas Dataset and Cell Type Annotation using Differentially Expressed Genes (DEGs) and Canonical Markers. (a) Top 10 DEGs per cluster identified by PCA. (b) Across-cluster average expression of cluster-specific DEGs. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. (d) Expression of human pancreatic canonical markers in PCA embeddings. The cell-type-specific markers considered include: GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells). UMAP visualizations show PCA cell embeddings, with cluster IDs assigned by K-means clustering. The overall clustering of cell types is limited compared to CoMem-DIPHW clustering, as evidenced by the overlapping beta cells, delta cells and PP cells in (d) and the lack of distinct high expression patterns of the DEGs in (b).
@@
In addition to the human pancreas results presented here, we also analyzed mouse pancreas, human brain, and mouse brain datasets. For these datasets, we have included visualizations in the appendix that show the average expression of the top DEGs across clusters identified by different methods, which shows our proposed algorithm’s ability in effectively distinguishing cell clusters by their expression profiles. The supplements also contain tables listing the top DEGs for clusters identified by CoMem-DIPHW and PCA in each dataset for reference.

### v4

@@ III.5 Results on the scMixology Benchmark Datasets
Dataset / Evaluation Measures GreedyModularity Louvain Infomap Leiden PCA t-SNE Node2Vec UMAP
scMixology 3 class
ARI $0.5836$ $\pm$ 0.000 $0.9933$ $\pm$ 0.000 $0.2327$ $\pm$ 0.000 $0.9933$ $\pm$ 0.000 $0.9900$ $\pm$ 0.000 $0.9898$ $\pm$ 0.000 $0.9913$ $\pm$ $0.0037$ $0.9898$ $\pm$ 0.000
NMI $0.7193$ $\pm$ 0.000 $0.9864$ $\pm$ 0.000 $0.3843$ $\pm$ 0.000 $0.9864$ $\pm$ 0.000 $0.9797$ $\pm$ 0.000 $0.9810$ $\pm$ 0.000 $0.9829$ $\pm$ $0.0075$ $0.9810$ $\pm$ 0.000
ACC $0.6929$ $\pm$ 0.000 $0.9978$ $\pm$ 0.000 $0.3370$ $\pm$ 0.000 $0.9978$ $\pm$ 0.000 $0.996$ $\pm$ 0.000 $0.9967$ $\pm$ 0.000 $0.9971$ $\pm$ $0.0013$ $0.9967$ $\pm$ 0.000
AMI $0.7189$ $\pm$ 0.000 $0.9864$ $\pm$ 0.000 $0.2335$ $\pm$ 0.000 $0.9864$ $\pm$ 0.000 $0.9796$ $\pm$ 0.000 $0.9810$ $\pm$ 0.000 $0.9829$ $\pm$ $0.0075$ $0.9810$ $\pm$ 0.000
F1 $0.5868$ $\pm$ 0.000 $0.9978$ $\pm$ 0.000 $0.4588$ $\pm$ 0.000 $0.9978$ $\pm$ 0.000 $0.9967$ $\pm$ 0.000 $0.9967$ $\pm$ 0.000 $0.9971$ $\pm$ $0.0013$ $0.9967$ $\pm$ 0.000
scMixology 5 class
ARI $0.6645$ $\pm$ 0.000 $0.6746$ $\pm$ 0.000 $0.3002$ $\pm$ 0.0008 $0.6746$ $\pm$ 0.000 $0.9873$ $\pm$ 0.000 $0.9861$ $\pm$ 0.000 $0.9451$ $\pm$ $0.0073$ $0.9838$ $\pm$ 0.000
NMI $0.7852$ $\pm$ 0.000 $0.7953$ $\pm$ 0.000 $0.4355$ $\pm$ $0.0003$ $0.7917$ $\pm$ 0.000 $0.9747$ $\pm$ 0.000 $0.9739$ $\pm$ 0.000 $0.9150$ $\pm$ $0.0086$ $0.9707$ $\pm$ 0.000
ACC $0.7325$ $\pm$ 0.000 $0.7361$ $\pm$ 0.000 $0.4150$ $\pm$ $0.0018$ $0.7358$ $\pm$ 0.000 $0.9941$ $\pm$ 0.000 $0.9939$ $\pm$ 0.000 $0.9621$ $\pm$ $0.0059$ $0.9929$ $\pm$ 0.000
AMI $0.7851$ $\pm$ 0.000 $0.7951$ $\pm$ 0.000 $0.2735$ $\pm$ $0.0003$ $0.7916$ $\pm$ 0.000 $0.9747$ $\pm$ 0.000 $0.9738$ $\pm$ 0.000 $0.9149$ $\pm$ $0.0086$ $0.9707$ $\pm$ 0.000
F1 $0.6556$ $\pm$ 0.000 $0.6576$ $\pm$ 0.000 $0.5250$ $\pm$ $0.0028$ $0.6571$ $\pm$ 0.000 $0.9941$ $\pm$ 0.000 $0.9939$ $\pm$ 0.000 $0.9618$ $\pm$ $0.0062$ $0.9929$ $\pm$ 0.000
Advanced ScRNA-seq and Hypergraph-Based Methods
Dataset / Evaluation Measure EDVW graph-sc tsImpute CAKE scASDC DIPHW CoMem CoMem-DIPHW
scMixology 3 class
ARI $0.5960$ $\pm$ $0.0454$ $0.7020$ $\pm$ $0.2650$ 0.997 $\pm$ 0.000 $0.8318$ $\pm$ $0.0196$ $0.4227$ $\pm$ $0.4342$ $0.9900$ $\pm$ $0.0052$ $0.9912$ $\pm$ $0.0019$ $0.9933$ $\pm$ 0.000
NMI $0.6140$ $\pm$ $0.0266$ $0.7654$ $\pm$ $0.1883$ 0.993 $\pm$ 0.000 $0.8609$ $\pm$ $0.0205$ $0.4500$ $\pm$ $0.4397$ $0.9803$ $\pm$ $0.0097$ $0.9829$ $\pm$ $0.0033$ $0.9864$ $\pm$ 0.000
ACC $0.7996$ $\pm$ $0.0557$ $0.8612$ $\pm$ $0.1275$ 0.999 $\pm$ 0.000 $0.9958$ $\pm$ $0.0011$ $0.6435$ $\pm$ $0.2956$ $0.9967$ $\pm$ $0.0018$ $0.9971$ $\pm$ $0.0006$ $0.9978$ $\pm$ 0.000
AMI $0.6132$ $\pm$ $0.0266$ $0.7649$ $\pm$ $0.1887$ 0.993 $\pm$ 0.000 $0.8605$ $\pm$ $0.0206$ $0.4479$ $\pm$ $0.4418$ $0.9803$ $\pm$ $0.0097$ $0.9829$ $\pm$ $0.0033$ $0.9864$ $\pm$ 0.000
F1 $0.7991$ $\pm$ $0.0561$ $0.8513$ $\pm$ $0.1380$ 0.999 $\pm$ 0.000 $0.9957$ $\pm$ $0.0011$ $0.5477$ $\pm$ $0.3810$ $0.9967$ $\pm$ $0.0018$ $0.9971$ $\pm$ $0.0006$ $0.9978$ $\pm$ 0.000
scMixology 5 class
ARI $0.6572$ $\pm$ $0.0221$ $0.9884$ $\pm$ $0.0015$ $0.9892$ $\pm$ 0.000 $0.7333$ $\pm$ $0.0746$ $0.3919$ $\pm$ $0.3602$ $0.9440$ $\pm$ $0.0031$ $0.9879$ $\pm$ $0.0013$ 0.990 $\pm$ 0.002
NMI $0.7421$ $\pm$ $0.0197$ $0.9782$ $\pm$ $0.0019$ $0.9796$ $\pm$ 0.000 $0.8435$ $\pm$ $0.0217$ $0.4575$ $\pm$ $0.4160$ $0.9092$ $\pm$ $0.0043$ $0.9771$ $\pm$ $0.0026$ 0.980 $\pm$ 0.003
ACC $0.6985$ $\pm$ $0.0161$ $0.9950$ $\pm$ $0.0006$ $0.9954$ $\pm$ 0.000 $0.8568$ $\pm$ $0.0192$ $0.5553$ $\pm$ $0.2156$ $0.9628$ $\pm$ $0.0020$ $0.9946$ $\pm$ $0.0007$ 0.996 $\pm$ 0.001
AMI $0.7418$ $\pm$ $0.0197$ $0.9782$ $\pm$ $0.0019$ $0.9796$ $\pm$ 0.000 $0.8431$ $\pm$ $0.0218$ $0.4565$ $\pm$ $0.4169$ $0.9091$ $\pm$ $0.0043$ $0.9771$ $\pm$ $0.0026$ 0.980 $\pm$ 0.003
F1 $0.7113$ $\pm$ $0.0143$ $0.9950$ $\pm$ $0.0006$ $0.9949$ $\pm$ $0.000$ $0.8579$ $\pm$ $0.0175$ $0.3443$ $\pm$ $0.2246$ $0.9628$ $\pm$ $0.0018$ $0.9946$ $\pm$ $0.0007$ 0.996 $\pm$ 0.001
Table 4: Performance of Clustering Methods on the scMixology Benchmark Datasets [tian2019benchmarking] across Five Evaluation Measures (ARI, NMI, ACC, AMI, and F1). The scMixology benchmark datasets include a 3-class and a 5-class dataset. Results for each are shown in the top and bottom sections, respectively. Best values are in bold, and second best are underlined. Results are reported as mean $\pm$ standard deviation over five independent runs. On the 3-class dataset, tsImpute achieved the highest scores across all evaluation measures, with CoMem-DIPHW, Louvain, and Leiden tied for second. On the more challenging 5-class dataset, CoMem-DIPHW achieved the best performance across all evaluation measures, outperforming tsImpute, which ranked second.
III.6 Results on the Human Pancreas Dataset
@@
For datasets without ground-truth annotations (specifically, human pancreas, mouse pancreas, human brain, and mouse brain), we qualitatively evaluate clustering performance using differential expression (DE) analysis. We present detailed DE analysis results for the human pancreas dataset using CoMem-DIPHW, PCA, graph-sc, tsImpute, CAKE, and scASDC in this section and provide results for the human brain, mouse brain, and mouse pancreas in the Supplementary Sections S7 and S8.
We report the top 10 differentially expressed genes (DEGs) from each human pancreas cell cluster identified by CoMem-DIPHW (Fig. 7a), PCA (Fig. 8a), graph-sc (Fig. 9a), tsImpute (Fig. 10a), CAKE (Fig. 11a), and scASDC (Fig. 12a). To evaluate cluster purity, we examine the expression of these DEGs across clusters (Fig. 7b, Fig. 8b, Fig. 9b, Fig. 10b, Fig. 11b, Fig. 12b). An effective clustering method should yield DEGs that uniquely characterize each cluster while showing minimal expression in other clusters.
@@
Violin plots show that the DEGs identified by CoMem-DIPHW exhibit more cluster-specific expression patterns than those derived from PCA, which tend to show higher expression in multiple clusters. For example, the average expression of the DEGs from PCA Cluster 1 is higher in Clusters 2, 5, and 6. Similarly, the DEGs from Cluster 7 show higher average expression in Cluster 6. Among the comparison methods, CAKE and scASDC produced highly distinctive DEGs, with scASDC yielding clusters that showed the strongest intra-cluster cohesion and inter-cluster separability. Graph-sc and PCA show less cluster-specific profiles, with DEGs frequently expressed across multiple clusters.
To assess the biological relevance of the clusters, we cross-reference the cluster-specific DEGs (see Fig. 7a) with markers from the PanglaoDB cell type marker database (version 27 March 2020)[franzen2019panglaodb]. Cell-type assignment is based on the overlap between the DEGs of each cluster and the cell type-specific markers in the PanglaoDB marker database, with a match score calculated as the proportion of matched markers (Table 7c). CoMem-DIPHW (Fig. 7c) successfully annotates all nine clusters with biologically relevant pancreatic cell types. CAKE produces similarly meaningful annotations, while graph-sc has several clusters with no annotation or biologically implausible labels.
@@
To further validate the clustering results, we examine the expression of canonical pancreatic marker genes [muraro2016single], including GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells) in the embeddings (Fig. 7d, Fig. 8d, Fig. 9d, Fig. 10d, Fig. 11d, Fig. 12d). Each marker is distinctly expressed in the embedding generated by CoMem-DIPHW, as shown in Fig. 7d, confirming that CoMem-DIPHW embeddings effectively distinguish different cell types. CAKE and scASDC also produce well-separated embeddings. In contrast, PCA embeddings show poor separation. The results from CoMem-DIPHW align well with the cell-type annotation by the PanglaoDB database. The only discrepancy is in Cluster 4. Matching the DEGs with PanglaoDB cell type markers suggests Pancreatic stellate cells (PSCs) or Granulosa cells, while the COL1A1 and ESAM expression plots indicate mesenchymal or endothelial cells. However, this apparent discrepancy can be explained by the fact that PSCs can differentiate into cells with endothelial-like properties [cheng2021generation], and granulosa cells can exhibit mesenchymal-like behaviors [jozkowiak2020stemness].
@@
Figure 7: Clustering Performance of CoMem-DIPHW on the Human Pancreas Dataset and Cell Type Annotation Using Differentially Expressed Genes (DEGs) and Canonical Markers. (a) Top 10 DEGs per cluster identified by CoMem-DIPHW. For example, Cluster 0 (acinar cells) includes genes associated with digestive enzyme production, a primary function of pancreatic acinar cells, including CPA1, CPA2 [tamura2018mutations], PLA2G1B [hui2019group], CTRB1 [pinho2011adult], and PRSS1 [masson2013conservative]. Clusters 2 and 7 (alpha cells) include markers GLS, CRYBA2, FAP, LOXL4, MAFB, and RGS4 (cluster 2), and LOXL4, PAX6, RGS4, and CRYBA2 (cluster 7). Cluster 3 (ductal cells) includes SLC4A4, ANXA4, KRT19, CFTR, CLDN1, and ONECUT2. Cluster 4, likely mesenchymal or endothelial cells, aligns with 9 of the top 10 DEGs found by CoMem-DIPHW, including COL4A2, COL4A1, SPARC, EMP1, IGFBP4, HTRA1, PXDN, LAMA4, and NID1. Cluster 5 (delta cells) includes GAD2, PCSK1, LEPR, and CASR, while Cluster 6 (beta cells) includes MAFA, ERO1LB, PDX1, IAPP, SYT13, HADH, and ADCYAP1. Cell type assignments are based on canonical marker expression patterns shown in (d). (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high within-cluster expression of DEGs and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and PanglaoDB cell type-specific markers, with match scores computed based on the proportion of matched markers. (d) Expression of human pancreatic canonical markers in CoMem-DIPHW embeddings. UMAP visualizations show CoMem-DIPHW cell embeddings, with cluster IDs assigned by K-means clustering. Each subplot displays the expression of a canonical marker gene for a specific human pancreatic cell type. The cell-type-specific markers considered include GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells). Color intensity represents gene expression levels, with red hues indicating higher expression. Clusters showing high expression of cell-type-specific canonical markers are assigned to the corresponding cell types. Overall, clustering quality is demonstrated by (b) high within-cluster DEG expression, (c) consistency between PanglaoDB-based annotation and clustering, and (d) distinct separation of cell types.
@@
Figure 8: Clustering Performance of PCA on the Human Pancreas Dataset and Cell Type Annotation Using Differentially Expressed Genes (DEGs) and Canonical Markers. (a) Top 10 DEGs per cluster identified by PCA. (b) Average expression of cluster-specific DEGs across clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. (d) Expression of human pancreatic canonical markers in PCA embeddings. UMAP visualizations show PCA cell embeddings, with cluster IDs assigned by K-means clustering. Overall, the clustering of cell types is limited compared to CoMem-DIPHW clustering, as shown by the overlapping beta cells, delta cells, and $P$ cells in (d) and the absence of distinct high expression patterns of the DEGs in (b).
Figure 9: Clustering Performance of graph-sc on the Human Pancreas Dataset and Cell Type Annotation Using Differentially Expressed Genes (DEGs) and Canonical Markers. (a) Top 10 DEGs per cluster identified by graph-sc. (b) Across expression of cluster-specific DEGs across clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. Graph-sc clusters generally show moderate specificity with substantial cross-cluster overlap. (c) Cell type annotation using the PanglaoDB marker database. Four clusters (1, 4, 7, and 8) yield no annotation, indicating that their DEGs did not match known cell type markers. (d) Expression of human pancreatic canonical markers in graph-sc embeddings. UMAP visualizations show graph-sc cell embeddings, with cluster IDs assigned by K-means clustering. Overall, cell cluster separation is limited.
Figure 10: Clustering Performance of tsImpute on the Human Pancreas Dataset and Cell Type Annotation Using Differentially Expressed Genes (DEGs) and Canonical Markers. (a) Top 10 DEGs per cluster identified by tsImpute combined with PCA embeddings and K-means. (b) Average expression of cluster-specific DEGs across clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. tsImpute DEGs show high specificity in Clusters 2, 4, 5, 6, 7, 8, and 9. (c) Cell type annotation using the PanglaoDB marker database. Several annotations are biologically implausible for pancreas tissue, including parietal cells (Cluster 1), salivary mucous cells (Cluster 3), gamma delta T cells (Cluster 5), and pulmonary vascular smooth muscle cells (Cluster 8). These unexpected annotations suggest that the imputation process may introduce artifacts that confound cell type identification. (d) Expression of human pancreatic canonical markers in tsImpute combined with PCA embeddings. UMAP visualizations show tsImpute plus PCA cell embeddings, with cluster IDs assigned by K-means clustering. tsImpute improves cluster separation compared to PCA without imputation.
Figure 11: Clustering Performance of CAKE on the Human Pancreas Dataset and Cell Type Annotation Using Differentially Expressed Genes (DEGs) and Canonical Markers. (a) Top 10 DEGs per cluster identified by CAKE. (b) Average expression of cluster-specific DEGs across clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. CAKE DEGs show good cluster specificity. (c) Cell type annotation using the PanglaoDB marker database. Most CAKE clusters correspond to biologically relevant cell types: gamma/$P$ cells (Clusters 0 and 3), ductal cells (Clusters 1 and 5), pancreatic stellate cells (Cluster 2), acinar cells (Cluster 4), alpha cells (Cluster 7), and beta cells (Cluster 8). (d) Expression of human pancreatic canonical markers in CAKE embeddings. UMAP visualizations show CAKE cell embeddings, with cluster IDs assigned by K-means clustering. Marker genes show distinct concentration in the well-separated CAKE embeddings: GCG in Cluster 7, INS in Cluster 8, SST in Clusters 0 and 3, PRSS1 in Cluster 4, and KRT19 in Cluster 1.
Figure 12: Clustering Performance of scASDC on the Human Pancreas Dataset and Cell Type Annotation Using Differentially Expressed Genes (DEGs) and Canonical Markers. (a) Top 10 DEGs per cluster identified by scASDC. (b) Average expression of cluster-specific DEGs across clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. scASDC DEGs show good cluster specificity. (c) Cell type annotation using the PanglaoDB marker database. While most clusters are matched to relevant pancreatic cell types, some are annotated with unexpected cell types, including retinal progenitor cells (Cluster 0) and nuocytes (Cluster 7). Clusters 1 and 3 received no annotation. (d) Expression of human pancreatic canonical markers in scASDC embeddings. scASDC embeddings show the best internal cohesion and separability. The marker genes show distinct concentration in the well-separated scASDC embeddings. However, the cell type annotations suggest that the embeddings may not necessarily correspond to biologically meaningful clusters.
@@
In addition to the human pancreas results presented here, we also analyze mouse pancreas, human brain, and mouse brain datasets. For these datasets, we include visualizations in the supplementary material in Figures S12b through S29b that show the average expression of the top DEGs across clusters identified by different methods. These results show our proposed algorithm’s ability in effectively distinguishing cell clusters by their expression profiles. The supplementary material (Figures S12a through S29a) also contain tables listing the top DEGs for clusters identified by CoMem-DIPHW, PCA, graph-sc, tsImpute, CAKE, and scASDC in each dataset for reference.
To complement the qualitative biological validation presented above, we further assess clustering performance using quantitative cluster quality measures, including the Silhouette score, Calinski-Harabasz index, Davies-Bouldin index, and Coefficient of Variation. These evaluations are conducted in two settings: (1) directly in the learned embedding space and (2) by evaluating the same cluster assignments on the preprocessed expression data used as input to the embedding algorithms. We refer the reader to Supplementary Tables S2 and S3. Evaluation in the embedding space captures the geometric separability and cohesion of clusters, whereas evaluation on the preprocessed data assesses whether the resulting cluster assignments capture meaningful structure in the input gene expression data. As demonstrated by differential expression and cell-type annotation analyzes in the human pancreas data set, strong embedding-space cohesion and separation alone do not necessarily guarantee biological interpretability.